OpenStudy (anonymous):

what would be the solution set of |R-3|=10

jimthompson5910 (jim_thompson5910):

|R-3|=10 breaks down into R-3 = 10 or R-3 = -10 solve each equation for R

OpenStudy (anonymous):

so it would be 13,-7

jimthompson5910 (jim_thompson5910):

yes those are your two solutions

OpenStudy (anonymous):

thank you